An insurance company insured 2000 scooter drivers, 4000 car drivers & 6000 truck drivers. The probabilities of their accident are 0.1, 0.3, and 0.2 respectively. One of the insured persons meets with an accident. What is the probability that he is a scooter driver?

1
Expert's answer
2021-12-07T06:48:42-0500

Let "E_1" , "E_2"  and "E_3"  be the events of a driver being a scooter driver, car driver and truck driver respectively. Let A be the event that the person meets with an accident.

There are 2000 insured scooter drivers, 4000 insured car drivers and 6000 insured truck drivers.

Total number of insured vehicle drivers = 2000 + 4000 + 6000 = 12000

"\therefore P\left(E_{1}\right)=\frac{2000}{12000}=\frac{1}{6}, P\left(E_{2}\right)=\frac{4000}{12000}=\frac{1}{3}, P\left(E_{3}\right)=\frac{6000}{12000}=\frac{1}{2}"


Also, we have:

 "P\left(A \mid E_{1}\right)=0.01=\frac{1}{100}"

"\begin{aligned}

&P\left(A \mid E_{2}\right)=0.03=\frac{3}{100} \\

&P\left(A \mid E_{3}\right)=0.15=\frac{15}{100}

\end{aligned}"

 Now, the probability that the insured person who meets with an accident is a scooter driver is "P\left(E_{1} \mid A\right)."

Using Bayes' theorem, we obtain:

"P\left(E_{1} \mid A\right)=\frac{P\left(E_{1}\right) \times P\left(A \mid E_{1}\right)}{P\left(E_{1}\right) \times P\left(A \mid E_{1}\right)+P\left(E_{2}\right) \times P\left(A \mid E_{2}\right)+P\left(E_{3}\right) \times P\left(A \mid E_{2}\right)}"

"\begin{aligned}
&=\frac{\frac{1}{6} \times \frac{1}{100}}{\frac{1}{6} \times \frac{1}{100}+\frac{1}{3} \times \frac{3}{100}+\frac{1}{2} \times \frac{15}{100}} \\
&=\frac{\frac{1}{6}}{\frac{1}{6}+1+\frac{15}{2}} \\
&=\frac{1}{6} \times \frac{6}{52} \\
&=\frac{1}{52}
\end{aligned}"